OpenStudy (ijarritos):

Simplify 5^-1(3^-2) A) 45 B) 1/45 C) 15^-3 D) 15^2

OpenStudy (mertsj):

\[5^{-1}=\frac{1}{5}\] \[3^{-2}=\frac{1}{3^2}=\frac{1}{9}\]

OpenStudy (mertsj):

Multiply the two numbers.

OpenStudy (ijarritos):

mertsj multiply what numbers what is the answer

OpenStudy (mertsj):

What does the problem say to multiply. Did you read the problem?